OpenStudy (anonymous):

The ratio of the quantities of sugar and flour needed to bake a cake is 2:5. What is the quantity of sugar needed for a cake, if 750 grams of flour are used to bake it? will medal

OpenStudy (igreen):

Oh, we have to set up a proportion. \(\dfrac{2}{5} = \dfrac{x}{750}\) Multiply 750 to both sides: \(\dfrac{2 \times 750}{5} = x\) Simplify the left side, what's 2 * 750?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

2* 750 =1500

OpenStudy (anonymous):

but you cancel the 0 from the 1500 with 750 and you get 150??

OpenStudy (igreen):

Yes, 1500 is correct, so we have: \(\dfrac{1500}{5} = x\) Now divide 1500 / 5

OpenStudy (anonymous):

1500/5=300

OpenStudy (igreen):

Yes, that's our answer.

OpenStudy (anonymous):

thansk so much

OpenStudy (igreen):

Np

OpenStudy (igreen):

Oh, and \(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\)

OpenStudy (anonymous):

thsnks